what will it cost to to carpet a rectangular floor measuring 15 feet by 21 feet if the carpet costs 20.97 per square yard

Found 2 solutions by rfer, oberobic:
Answer by rfer(16322)

15*21=315
315/9=35 sq yds
35*20.97=$733.95

Answer by oberobic(2304)

area = 15 ft * 21 ft = 315 sq ft
but you need this in yards
1 sq yd = 3 ft * 3 ft = 9 sq ft
.
315 sq ft / 9 sq ft/yd = 35 sq yd
.
35 sq yd * $20.97 / sq yd = $733.9499999999999 or about $733.95
.
Done.